OpenStudy (anonymous):

0.12x-0.54=1.43x-0.76x-3.523

OpenStudy (anonymous):

Do you want help solving this or just the answer?

OpenStudy (anonymous):

help

OpenStudy (anonymous):

i know combine like terms

OpenStudy (anonymous):

Well, all you need to do is move the x terms to one side and have everything else on the other. Then you can factor out an x term and divide off the rest

OpenStudy (anonymous):

but it never comes out right

OpenStudy (anonymous):

Put all the X's on one side, and all the constants on the other. So do: -0.54+3.523 = 1.43x - 0.76x - 0.12x and try to go from there

OpenStudy (anonymous):

Once you combine all of those terms, then divide everything by the coefficient of X (number in front of x)

OpenStudy (anonymous):

but it never comes out right

OpenStudy (anonymous):

4.5

OpenStudy (anonymous):

but it never comes out right

OpenStudy (anonymous):

???

OpenStudy (anonymous):

you should be getting 5.4

OpenStudy (anonymous):

Here is a worked solution the your problem, compare it with your notes and see where the trouble lies. I was careful not to skip any steps: So we have -0.54+3.523 = 1.43x - 0.76x - 0.12x 2.983 = 1.43x - 0.76x - 0.12x 2.983 = 1.43x - 0.88x 2.983 = 0.55x 2.983 / 0.55 = 0.55x / 0.55 x = 5.423636363636...

OpenStudy (anonymous):

DS is faster than me :P \[0.12x-1.43x+0.76x=-3.523+0.54\] \[0.12x-1.43x+0.76x=-2.983\] \[x(0.12-1.43+0.76)=-2.983\] \[x(0.12-1.43+0.76)=-2.983\] \[x=(-2.983)/(0.12-1.43+0.76)\] \[x=(-2.983)/(-0.55)\] \[x=5.42\]

OpenStudy (anonymous):

but it never comes out right

OpenStudy (anonymous):

woah thats alot

OpenStudy (anonymous):

Maybe you are missing a negative sign somewhere